OpenStudy (anonymous):

The range of the function f(x)=x+5 is {7,9} what is the domain?

OpenStudy (danjs):

the range is 2 values, those will correspond to 2 points on the graph

OpenStudy (danjs):

range are the Y values, use 7 and 9 for Y and solve what X will be for each of those

OpenStudy (danjs):

f(x) = y

OpenStudy (danjs):

7 = x + 5 and 9 = x + 5 those 2 x values are your domain,

OpenStudy (anonymous):

Thank you!